OpenStudy (anonymous):

7/2xy - 5/2xy ...the answer is 1/xy .....but how do you figure it out step by step?

OpenStudy (anonymous):

these are 2 fractions having the same denominator,so you just have to subtract the numerators: 7/2xy - 5/2xy=(7-2)/(2xy)=2/(2xy)=1/(xy)

OpenStudy (anonymous):

thank you :)